OpenStudy (sgayton27):

simplify the power of i: i^-55

OpenStudy (sgayton27):

\[i ^{-55}\]

OpenStudy (anonymous):

OGH not these again @ranga

OpenStudy (anonymous):

Reaper help i have one more question for you help this one first

OpenStudy (anonymous):

mssg me a link

OpenStudy (amistre64):

i is a circular construction, it repeats every 4 times what is 55/4 ?

OpenStudy (amistre64):

it might be better asked what the remainder is when -55 is divided by 4, but those are semantics

OpenStudy (sgayton27):

3

OpenStudy (sgayton27):

is the remainder

OpenStudy (ranga):

i = i i^2 = -1 i^3 = -i i^4 = +1 i^5 = i (same as first line. starts repeating) remainder of 55/4 is 3. So it will be the third line: -i i^(-55) = 1 / i^(+55) = 1 / -i = -i / (-i * -i) = -i / +1 = -i

OpenStudy (sgayton27):

i^-55 is the same as \[\frac{ 1 }{ i ^{55} }\]

OpenStudy (ranga):

Yes. See the above answer.

OpenStudy (amistre64):

i^(-3) not= i^(3)

OpenStudy (sgayton27):

well with the remainder being 3 we get \[\frac{ 1 }{ i ^{3} }\] which is the same as \[\frac{ 1 }{ 1^{2}*i }\] which equals \[\frac{ 1 }{ (-1)*i }\] so the answer is \[\frac{ 1 }{ -i }\]

OpenStudy (amistre64):

its the same as i^1 v -3 -2 -1 0 1 2 3 4 5 6 7

OpenStudy (amistre64):

the issue with negative exponents is that they are just shy of being expressed as postive exponents. -3 + 4 = 1 converts it to a usual i^+

OpenStudy (ranga):

Oh, I didn't realize the typo in my last line: i^(-55) = 1 / i^(+55) = 1 / -i = i / (-i * i) = i / +1 = i

OpenStudy (ranga):

@sgayton27 Yes, 1 / -i except you don't leave an i in the denominator. You rationalize the denominator (meaning you make the denominator a rational number) by multiplying top and bottom by i:\[\frac{ 1 }{ -i } = \frac{ 1 * i }{ -i * i } = \frac{ i }{ -i^2 } = \frac{ i }{ -(-1) } = \frac{ i }{ +1 } = i\]